# Experimental verification of the quantum nature of a neural network

### v4

@@ I Introduction
A (cvasi)-classical neural network has certain properties that could be assimilated as quantum to some extent. While we would not propagate a quantum state through a series of quantum devices forming a hard to maintain network, we would instead use a set of classical neurons that, acting together, would be able to produce a state that could be classified as quantum for all practical purposes. Let us consider a neural network, made out of neurons firing according to classical rules. Let us also consider several layers of such neurons, as can be imagined in a deep learning network. The parameters of the neurons are bound together in a series of linear combinations in which the activation function plays the role of a non-linear contributor to the adaptability of the network. The relation among the neurons however is always linear. A signal is transmitted to the network as an input, which is passed then through the network and a gradient descent backpropagation method is used to train the linearly connected neurons forming the network to the desired structure. After defining a loss function, backpropagation is required to transmit the end-point error back to the input node and to optimise the parameters in the process. This process allows the network to access, via its non-linear activation function and particularly through its loss function, the non-local structure of the features it explores. This is also why neural networks are so good at classifying non-local classification problems and are finally used as approximations for highly entangled quantum states. But are classical neural networks just approximations of quantum systems, or are they, in some sense, quantum themselves, not through the interactions that make them function, but through their global structure and their backpropagation and loss function features (mainly their dynamics maps)? Usually, in a neural network, the optimised result is not easily associated with the input, which is why the artificial neural networks have the well known problem of inferring causal connections. This may not be surprising given that the end-result is obtained through a process of extensive linear superposition of learned data and then optimised via a loss function that introduces the learned data to the global structure of the problem it is asked to solve. This is what an entangling gate would do. A Hadamard gate would create a linear superposition of the initial state. This is what occurs to our initial data via the network linear superposition. A CNOT gate would combine two states, switching the result of one when the other is in a specified state. This is generally the result of our loss functions. The only difference is that we are not a-priori acting on quantum states but on suitably represented initial learning data. It all depends on what meaning we give to that data and how we interpret the output then, to make it a quantum analogue or a simple classical superposition.

### v5

@@ I Introduction
Let us consider a neural network, made out of neurons firing according to classical rules. Let us also consider several layers of such neurons, as can be imagined in a deep learning network. The parameters of the neurons are bound together in a series of linear combinations in which the activation function plays the role of a non-linear contributor to the adaptability of the network. The relation among the neurons however is always linear. A signal is transmitted to the network as an input, which is passed then through the network and a gradient descent backpropagation method is used to train the linearly connected neurons forming the network to the desired structure. After defining a loss function, backpropagation is required to transmit the end-point error back to the input node and to optimise the parameters in the process. This process allows the network to access, via its non-linear activation function and particularly through its loss function, the non-local structure of the features it explores. This is also why neural networks are so good at classifying non-local classification problems and are finally used as approximations for highly entangled quantum states. But are classical neural networks just approximations of quantum systems, or are they, in some sense, quantum themselves, not through the interactions that make them function, but through their global structure and their backpropagation and loss function features (mainly their dynamics maps)? Usually, in a neural network, the optimised result is not easily associated with the input, which is why artificial neural networks have the well known problem of inferring causal connections. This may not be surprising given that the end-result is obtained through a process of extensive linear superposition of learned data and then optimised via a loss function that introduces the learned data to the global structure of the problem it is asked to solve. This is what an entangling gate would do.
A Hadamard gate would create a linear superposition of the initial state. This is what occurs to our initial data via the network linear superposition. A CNOT gate would combine two states, switching the result of one when the other is in a specified state. This is generally the result of our loss functions. The only difference is that we are not a-priori acting on quantum states but on suitably represented initial learning data. It all depends on what meaning we give to that data and how we interpret the output then, to make it a quantum analogue or a simple classical superposition.
In nature, biological neural networks work through the process of accepting signals in the form of collections of molecular ions that reach the synapse passing through the pre-synaptic region. They are not unique signals, but statistical distributions of such molecular ions that present strong correlations with each other due to the mechanisms that produced them in the pre-synaptic region. Therefore the real situation in nature is much more complicated than what the current simplified models present us. Treating the input as a well define classical signal is certainly an over-simplification. The signal in fact results as an emergent property of a distribution of molecular ions. Whether the distribution has or has not quantum correlation is a question not completely answered. In particular, the whole process of transmitting the distribution of molecular ions in the synaptic region, receiving it by the receptors of the neuron, and then triggering a series of mechanisms inside the neuron could be a quasi-classical process, but how exactly the quasi-classical dynamics is obtained as a limit of quantum phenomenon is not understood as for now.
As I discussed in the previous chapter, taking the classical limit $\hbar\rightarrow 0$ is not trivial and in most cases the non-linear classical dynamics in a neural network would create quite interesting problems. In order to describe such processes, I bring in an example of working with the Wigner-Weyl representation. The main goal of this approach is to find a structure that comes closest to a quantum distribution over the phase space. This is of course technically impossible due to the non-commutativity of momentum and position in quantum mechanics. If we want to calculate expectation values, which is the case for example in the applications of Ehrenfest’s theorem, we need some form of distributions. This can be achieved by considering Wigner distributions and Weyl transformed operators. Indeed, given a density matrix
$\hat{\rho}=\left|\psi\right\rangle\left\langle\psi\right|$ (1)
we combine the process of having a Weyl transformed operator
[…]
and the formation of a Wigner distribution starting from the expectation value of the density matrix
$\left\langle x\right|\hat{\rho}\left|x^{\prime}\right\rangle=\psi(x)\psi^{*}(x^{\prime})$ (3)
in the form
[…]
Defining the trace of two operators as
$Tr[\hat{A}\hat{B}=\frac{1}{h}\int\int\tilde{A}(x,p)\tilde{B}(x,p)dxdp$ (5)
we obtain for the expectation value of an operator
[…]
If we want to analyse the dynamics we look at the partial time derivatives of the Wigner distribution
[…]
By looking at the Schrodinger equation
$\frac{\partial\psi(x,t)}{\partial t}=-\frac{\hbar}{2im}\frac{\partial^{2}\psi(x,t)}{\partial x^{2}}+\frac{1}{i\hbar}U(x)\psi(x,t)$ (8)
we can split the time evolution of the Wigner distribution as in
$\frac{\partial W}{\partial t}=\frac{W_{T}}{\partial t}+\frac{\partial W_{U}}{\partial t}$ (9)
The kinetic part we are treating easily as
[…]
which can be written as
$\frac{\partial W_{T}}{\partial t}=-\frac{p}{m}\frac{\partial W(x,p)}{\partial x}$ (11)
and the potential part which can be written as
[…]
which by means of a series expansion can be written as
$\frac{\partial W_{U}}{\partial t}=\sum_{s=0}(-\hbar^{2})^{s}\frac{1}{(2s+1)!}(\frac{1}{2})^{2s}\frac{\partial^{2s+1}U(x)}{\partial x^{2s+1}}(\frac{\partial}{\partial p})^{2s+1}W(x,p)$ (13)
we see however that taking the $\hbar\rightarrow 0$ limit here is not trivial as we also have a $(2s+1)$ derivative of $W(x,p)$ which would always bring us more $\hbar$ to the denominator, than what we could eliminate and hence the $\hbar\rightarrow 0$ limit would be ill-defined. Usually to avoid this problem, the transition from quantum to classical is taken by introducing mixed states. Indeed, classical distributions are linear in phase space, while Wigner functions are not. We can write linear superpositions of wavefunctions $\psi=\psi_{\alpha}+\psi_{\beta}$ but from that it doesn’t follow the superposition of the associated Wigner functions, therefore we cannot conclude that $W_{\psi}=W_{\alpha}+W_{\beta}$. If we introduce mixed states, however
$W(x,p)=\tilde{\rho}/h=\sum_{j}P_{j}W_{j}(x,p)$ (14)
we recover the linearity as is the case for classical phase space distributions, but indeed, we use mixed states. The advantage is that, while indeed we use mixed states, we can see easier how the quantum terms appear and in particular we can safely take the $\hbar\rightarrow 0$ limit. For the sake of an example, I will use as basic model the harmonic oscillator ground state. In the case we consider pure states, we would have
$\psi=A[\psi_{0}(x-b)+\psi_{0}(x+b)]$ (15)
and we use the ground state of the harmonic oscillator
$\psi_{0}(x)=\frac{1}{\pi^{1/4}\sqrt{a}}e^{-x^{2}/(2a^{2})}$ (16)
leading to the Wigner distribution
$W(x,p)=\frac{1}{h(1+e^{-b^{2}/a^{2}})}e^{-(ap)^{2}/\hbar^{2}}[e^{-(x-b)^{2}/a^{2}}+e^{-(x+b)^{2}/a^{2}}+2e^{-x^{2}/a^{2}}cos92bp/\hbar)]$ (17)
In this case we would have a superposition of quantum states. If we work however with mixed states we write a linear combination of classical states and therefore we obtain
$W(x,p)=\frac{1}{2}[W_{0}(x-b,p)+W(x+b,p)]$ (18)
where
$W_{0}(x,p)=\frac{2}{h}exp(-a^{2}p^{2}/\hbar-x^{2}/a^{2})$ (19)
and therefore
$W(x,p)=\frac{1}{h}e^{-a^{2}p^{2}/\hbar^{2}}[e^{-(x-b)^{2}/a^{2}}+e^{-(x+b)^{2}/a^{2}}]$ (20)
Applying now the partial derivative $(\frac{\partial}{\partial p})^{2s+1}$ on the Wigner functions of mixed states
[…]
defined by means of a probability density
$P(p_{0})=\frac{1}{c\sqrt{\pi}}e^{-p_{0}^{2}/c^{2}}$ (22)
results in
$W(x,p)=\frac{1}{\pi\sqrt{a^{2}c^{2}+\hbar^{2}}}e^{-x^{2}/a^{2}}e^{p^{2}/(c^{2}+\hbar^{2}/a^{2})}$ (23)
In general we couldn’t fix $p$ or $x$ because of the uncertainty relations of quantum mechanics that links them through $\hbar$. Indeed the dynamics can reverse the places of x and p (we can obtain squeezed states, etc.) therefore fixing one is not realistic. If we fix the width on $x$, then the width on $p$ cannot be arbitrarily small. However, once we shift to mixed states, the width of other $p$-states can be arbitrary (large or small) and therefore we can safely take the limit $\hbar\rightarrow 0$. In this way in the example of a double coherent pure state, taking the classical limit by means of mixed states, we obtain
$W(x,p)=\frac{1}{2\pi ac(1+e^{-(b/a)^{2}})}e^{-p^{2}/c^{2}}[e^{-(x-b)^{2}/a^{2}}+e^{-(x+b)^{2}/a^{2}}+2e^{-x^{2}/a^{2}}e^{-b^{2}/a^{2}}]$ (24)
where the last term expresses the level of suppression of quantum properties in a continuous manner. However, non-linear classical dynamics in this approach to transitioning to the classical limit may result in very large higher order derivatives. Combined with our $\hbar\rightarrow 0$ limit, this may indeed amplify quantum effects by means of classical non-linear dynamics. As neural networks do implement non-linear classical dynamics, it seems that controlling the non-linearity parameters of a quasi-classical neural network may induce or amplify quantum effects.